# 11footnotetext: Equal Contributions

### v2

@@ Abstract
We address key challenges in Dataset Aggregation (DAgger) for real-world contact-rich manipulation: how to collect informative human correction data and how to effectively update policies with this new data. We introduce Compliant Residual DAgger (CR-DAgger), which contains two novel components: 1) a Compliant Intervention Interface that leverages compliance control, allowing humans to provide gentle, accurate delta action corrections without interrupting the ongoing robot policy execution; and 2) a Compliant Residual Policy formulation that learns from human corrections while incorporating force feedback and force control. Our system significantly enhances performance on precise contact-rich manipulation tasks using minimal correction data, improving base policy success rates by over 50% on two challenging tasks (book flipping and belt assembly) while outperforming both retraining-from-scratch and finetuning approaches. Through extensive real-world experiments, we provide practical guidance for implementing effective DAgger in real-world robot learning tasks. Result videos are available at: https://compliant-residual-dagger.github.io/
@@ 1 Introduction
Learning from human demonstrations has seen many recent successes in real-world robotic tasks [8, 19, 37, 42]. However, to obtain a successful policy, human demonstrators often have to repeatedly deploy a policy and observe its failure cases, then collect more data to update the policy until it succeeds. This process is broadly referred to as Dataset Aggregation (DAgger) [33, 21]. However, doing DAgger effectively for real-world robotic problems still faces the following challenges:
How to collect informative human correction data? DAgger is most effective when the correction data is within the original policy’s induced state distribution [33]. In practice, the common approach is either (1) collecting offline demonstrations that cover the policy’s typical failure scenarios [7], or (2) human taking over robot control during policy deployment [34, 30]. However, in both cases, it is difficult for human demonstrators to gain the intuition for how to execute the corrective actions without deviating excessively from the original distribution. Human taking over additionally introduces force discontinuity when they do not instantly reproduce the exact same robot force. This is partially due to the lack of effective correction interfaces that support precise and instantaneous intervention.
How to effectively update the policy with new data? Prior methods for improving a pretrained policy with additional data include (1) retraining the policy from scratch with the aggregated dataset [21], which can be computationally expensive; (2) finetuning the policy with only the additional data [37, 15, 6], which is sensitive to the quality of the new data [45], and (3) training a residual policy separately on top of the pretrained policy, which is typically done with Reinforcement Learning  [1, 45] or Imitation Learning [4], both require a large number of samples.
@@
  • Compliant Intervention Interface. We propose an on-policy correction system based on kinesthetic teaching to collect delta action without interrupting the current robot policy. Leveraging compliance control, the interface lets humans directly feel the magnitude of their instantaneous correction, so they can provide gentle adjustments. Unlike take-over corrections that may cause force discontinuity, our design allows smooth transition between correction/no correction mode, while maintaining distributional consistency with the original policy.
@@
Together, our system significantly improves the success rate of precise contact-rich robot manipulation tasks using a small amount of additional data. We demonstrate the efficacy of our method on two challenging tasks with long horizons and sequences of contacts: book flipping and belt assembly. We improve over the base policy success rate by over 50% using less than 50 intervention episodes, while also outperforming retrain-from-scratch and finetuning under the same data budgets. In summary, our contributions are:
  • A Compliant Intervention Interface, a system that allows human to provide accurate, gentle, and smooth corrections in both position and force to a running robot policy without interrupting it.
@@
  • A practical guidance for efficient DAgger based on extensive real-world studies for critical but often overlooked design choices such as batch size and sampling strategy. Our hardware design, training data, and policy code will be made publicly available.
@@ 2 Related Work
Human-in-the-Loop Corrections for Robot Policy Learning. The original DAgger work [33] requires the demonstrator to directly label actions generated by the policy. In robotics, a practical variation is to let the human take over the robot control and provide correct action directly [21]. Such human correction motion can be recorded with spacemouse [6, 27], joystick [34], smartphone [30], or arm-based teleoperation system [16, 17, 37]. We instead proposes a novel kinesthetic teaching system with compliance controller that allows the demonstrator to apply delta corrections while the robot policy is still running, and additionally records force feedback. Our results show that both the delta correction data and the force data are crucial to the success of the learned policy.
Improving Pretrained Robot Policies with New Data. The most direct approach to improving a pretrained policy with new correction data is to retrain the policy on the aggregated dataset, combining prior demonstrations with new feedback [30, 34]. Alternatively, Reinforcement Learning (RL) offers a framework to incorporate both offline and online data, either by warm-starting replay buffers [29, 2] or by using offline data to guide online fine-tuning [43, 38]. When policies are trained on large-scale human demonstration datasets [5, 35, 23, 32, 22, 39, 40], retraining becomes impractical, especially when the original data is inaccessible. In such cases, fine-tuning with only the new data is a common solution, using either imitation learning [27, 15, 37] or RL [31, 6]. Another line of work introduces an additional residual model on top of the original policy. These residual policies can be trained with RL in simulation [45, 1, 13], but suffers from sim-to-real challenges. Training residual policy in the real world usually requires a large number of samples [4, 20], intermediate scene representation [12], or consistent visual observations between training and testing [11, 14], making the approach hard to adopt in practice. In this work, we introduce a practical data collection system and an efficient residual policy learning algorithm for long-horizon, contact-rich manipulation tasks. Our approach requires only a small amount of real-world correction data and supports integration of additional sensory modalities not present in the original model, leading to improved policy performance.
@@ 3.1 Compliant Intervention Interface
Correction data is collected by human demonstrators to rectify policy failures. Unlike initial demonstrations that establish baseline behaviors, correction data specifically targets failure modes observed during policy deployment. Correction data is most effective when it corrects failures in policy-induced state distributions [33]. The interface through which these corrections are collected significantly impacts the quality of correction data, which should be intuitive for demonstrators, capture critical corrective information at precise moments of failures, and help correction data maintain distributional consistency with base policy outputs.
There are two types of correction collection methods: Off-policy correction is when humans observe failures of the base policy during deployment, and then recollect extra offline demonstrations to address failure cases. This approach is most commonly used for improving Behavior Cloning policy performance due to its simplicity - it requires no additional infrastructure beyond the original data collection setup. However, the resulting demonstrations may fail to cover all the failure cases or deviate from the policy’s original distribution. We focus on on-policy correction instead, where humans can monitor policy execution and intervene on the spot when failures occur or are anticipated. This approach allows humans to provide corrections more targeted to the base policy’s failure cases. However, challenges still exist for an intervention system:
@@
  • Distribution shift. The human intervention may still introduce significant distribution shifts as the motion deviates too much from the original policy distribution. Additionally, the non-smooth transition above could bring in disturbances and add to the distribution shift.
  • Indirect correction brings errors. Correction is commonly implemented via teleoperation interfaces such as spacemouse or joysticks [15, 6]. With spatial mismatch and teleoperation latency, it is hard for the demonstrator to instantly provide accurate corrections upon intervention starts without going through a short adjustment period.
@@
  • Delta correction instead of take-over correction. Unlike take-over correction, where the demonstrator has no idea of the policy’s original intention once taking over, we propose a novel on-policy delta correction method: we let the robot policy executes continuously while the human applies forces to the robot with a handle mounted on the end effector, resulting in delta actions on top of the policy action. The human demonstrator can always sense the policy’s intention through haptic feedback, and easily control the magnitude of intervention by the amount of force applied to the handle. As a result, delta correction ensures smooth intervention data and limits the human from providing very large corrections that can easily lead to out-of-distribution states. The approach is also intuitive as the human can directly move the robot towards desired correction directions.
[…]
@@
[…]
@@ 3.2 Compliant Residual Policy
The residual policy directly uses the base policy’s frozen image encoder to extract an image embedding, a temporal convolution network [36] to encode the force vectors, followed by fully-connected layers to decode actions.
Advantages. This formulation provides the following advantages:
  • Sample-efficient learning. The residual policy’s network is light-weight ($\sim$2MB trainable weights) and only requires a small amount of correction data to train ($\sim$50 demonstrations).
@@
Training strategy. In prior work, a residual policy is trained either in simulation with RL [1, 45] to give it sufficient coverage of the input distribution, or in the real world with pre-collected behavior cloning data [29]. In this work, we train the Compliant Residual Policy completely on the small amount of new real-world correction data with the following strategies:
@@
  • Prioritize correction data over no-correction (zero residual action) data. Similar to [27], we alter the sample frequency of intervention data during training based on whether they have human correction or not. Specifically, since the moment of correction start indicates where the current policy performs badly followed by immediate action to fix it, we sample data more frequently for a short period immediately after correction starts. Our real-world ablations (§ 4.5) demonstrate that our training strategies improve the quality of the residual policy and the overall success rate.
@@ 4 Evaluation
For each task, we train a diffusion policy [7] with 150 demonstrations as the base policy. We first deploy the base policy and observe its performance and failure modes. Next, from the same base policy, we collect 50 correction episodes with each data collection method. Then, we update the policy using each network updating method and training procedure. Finally, we deploy the updated policies and evaluate their performance under the same test cases. Each data & policy combination is evaluated on each task for 20 trials. Details of tasks and comparisons are described below.
@@ 4.1 Contact-Rich Manipulation Tasks
Figure 5: Belt Assembly Task. (a) Policy rollout of [Compliant Residual] policy trained with [On-Policy Delta] data, demonstrating accurate force-position coordination and adaptation. (b) Different test scenarios. (c) Typical failure cases of the base policy: getting stuck on base next to the small pulley; blocked by the the button on the large pulley; and missing the slot by going too high above the pulley.
Book Flipping: As shown in Fig. 4 (a), this task is to flip up books on a shelf. Starting with one or more books lying flat on the shelf, the robot first insert fingers below the book, then rotate the book up and push them firmly against the shelf to let them stand on their own.
This task is challenging as it involves rich use of physical contacts and forceful strategies [18]. A position-only strategy always fails immediately by triggering large forces, so we execute all policies through the same admittance controller. The task success requires high precision in both motion and force to accurately align the fingers with the gap upon insertion, and to provide enough force to rotate heavy books and make the books stand firmly.
@@
[…]
The task is challenging as it requires both position and force accuracy throughout the process. Specifically, the belt is thin and soft so the initial alignments onto the pulleys are visually ambiguous. Also, since the belt is not stretchable, there is more resistance force and less position tolerance as the belt approaches the second pulley, requiring a policy with good force-position coordination and adaptation. We ran 20 rollouts across 4 different initial board positions for all methods (Fig. 5 (b)). We use the same test cases and initial robot configuration for all evaluations.
@@ 4.2 Base Policy and its Failure Modes
We trained a diffusion policy [7] that takes in past images from a wrist-mounted camera and robot proprioception observations, and predicts a future position-based action trajectory. We collected 150 successful demonstrations on both tasks, with five different books for the book flipping task, and randomized initial robot and object configurations for both tasks.
@@
The belt assembly base policy achieves a 20% success rate with the following common failure cases (Fig. 5 (c)): (1) Missed small pulley assembly: the robot moves with an improper height, causing the belt to get stuck at the base of the board next to the small pulley. (2) Missed big pulley assembly: the fingertip goes too high or too low, causing the belt to miss the slot on the big pulley.

### v3

@@ Abstract
We address key challenges in Dataset Aggregation (DAgger) for real-world contact-rich manipulation: how to collect informative human correction data and how to effectively update policies with this new data. We introduce Compliant Residual DAgger (CR-DAgger), which contains two novel components: 1) a Compliant Intervention Interface that leverages compliance control, allowing humans to provide gentle, accurate delta action corrections without interrupting the ongoing robot policy execution; and 2) a Compliant Residual Policy formulation that learns from human corrections while incorporating force feedback and force control. Our system significantly enhances performance on precise contact-rich manipulation tasks using minimal correction data, improving base policy success rates by over 60% on two challenging tasks (book flipping and belt assembly) while outperforming both retraining-from-scratch and finetuning approaches. Through extensive real-world experiments, we provide practical guidance for implementing effective DAgger in real-world robot learning tasks. Result videos are available at: https://compliant-residual-dagger.github.io/
@@ 1 Introduction
Learning from human demonstrations has seen many recent successes in real-world robotic tasks [chi2024universal, hou2024adaptive, wu2025robocopilot, xu2025robopanoptes, xiong2025vision]. However, to obtain a successful policy, human demonstrators often have to repeatedly deploy a policy and observe its failure cases, then collect more data to update the policy until it succeeds. This process is broadly referred to as Dataset Aggregation (DAgger) [ross2011reduction, kelly2019hg]. However, doing DAgger effectively for real-world robotic problems still faces the following challenges:
How to collect informative human correction data? DAgger is most effective when the correction data is within the original policy’s induced state-action distribution [ross2011reduction]. In practice, the common approach is either (1) collecting offline demonstrations that cover the policy’s typical failure scenarios [chi2023diffusion], or (2) human taking over robot control during policy deployment [spencer2020learning, mandlekar2020human]. However, in both cases, the human demonstrator has no access to the original policy’s behavior and may deviate excessively from it. Human taking over additionally introduces force discontinuity when they do not instantly reproduce the exact same robot force. This is partially due to the lack of effective correction interfaces that support precise and instantaneous intervention.
How to effectively update the policy with new data? Prior methods for improving a pretrained policy with additional data include (1) retraining the policy from scratch with the aggregated dataset [kelly2019hg], which can be computationally expensive; (2) finetuning the policy with only the additional data [wu2025robocopilot, he2025uncertainty, chen2025conrft], which is sensitive to the quality of the new data [yuan2024policy], and (3) training a residual policy separately on top of the pretrained policy, which is typically done with Reinforcement Learning  [ankile2024imitation, yuan2024policy] or Imitation Learning [bharadhwaj2024track2act], both require a large number of samples.
@@
  • Compliant Intervention Interface. We propose an on-policy correction system based on kinesthetic teaching to collect delta action without interrupting the current robot policy. Leveraging compliance control, the interface lets humans directly apply force to the robot and feel the magnitude of their instantaneous correction. Unlike take-over corrections, our design allows smooth transition between correction/no correction mode, while providing direct control of correction magnitudes.
@@
Together, our system significantly improves the success rate of precise contact-rich robot manipulation tasks using a small amount of additional data. We demonstrate the efficacy of our method on two challenging tasks involving long horizons and sequences of contacts: book flipping and belt assembly. We improve over the base policy success rate by over 60%, while also outperforming retrain-from-scratch and finetuning under the same data budgets. In summary, our contributions are:
  • A Compliant Intervention Interface, a system that allows humans to provide accurate, gentle, and smooth corrections in both position and force to a running robot policy without interrupting it.
@@
  • A practical guide for efficient DAgger based on extensive real-world studies for critical but often overlooked design choices, such as batch size and sampling strategy. Our hardware design, training data, and policy code can be found here.
@@ 2 Related Work
Human-in-the-Loop Corrections for Robot Policy Learning. The original DAgger work [ross2011reduction] requires the demonstrator to directly label actions generated by the policy. In robotics, a practical variation is to let the human take over the robot control and provide correct action directly [kelly2019hg]. Such human correction motion can be recorded with spacemouse [chen2025conrft, liu2022robot], joystick [spencer2020learning], smartphone [mandlekar2020human], or arm-based teleoperation system [hoque2021thriftydagger, hoque2021lazydagger, wu2025robocopilot]. We instead proposes a novel kinesthetic teaching system with compliance controller that allows the demonstrator to apply delta corrections while the robot policy is still running, and additionally records force feedback. Our results show that both the delta correction data and the force data are crucial to the success of the learned policy.
Improving Pretrained Robot Policies with New Data. The most direct approach to improving a pretrained policy with new correction data is to retrain the policy on the aggregated dataset, combining prior demonstrations with new feedback [mandlekar2020human, spencer2020learning]. Alternatively, Reinforcement Learning (RL) offers a framework to incorporate both offline and online data, either by warm-starting replay buffers [luo2024precise, ball2023efficient] or by using offline data to guide online fine-tuning [yin2025rapidly, xu2022aspire]. When policies are trained on large-scale human demonstration datasets [black2024pi_0, team2024octo, kim2024openvla, o2024open, khazatsky2024droid, xu2023xskill, xu2024flow], retraining becomes impractical, especially when the original data is inaccessible. In such cases, fine-tuning with only the new data is a common solution, using either imitation learning [liu2022robot, he2025uncertainty, wu2025robocopilot] or RL [mark2024policy, chen2025conrft]. Another line of work introduces an additional residual model on top of the original policy. These residual policies can be trained with RL in simulation [yuan2024policy, ankile2024imitation, haldar2023watch], but suffers from sim-to-real challenges. Training residual policy in the real world usually requires a large number of samples [bharadhwaj2024track2act, johannink2019residual], intermediate scene representation [guzey2024bridging], or consistent visual observations between training and testing [guzey2024see, haldar2023teach], making the approach hard to adopt in practice. In this work, we introduce a practical data collection system and an efficient residual policy learning algorithm for long-horizon, contact-rich manipulation tasks. Our approach requires only a small amount of real-world correction data and supports integration of additional sensory modalities not present in the original model, leading to improved policy performance.
@@ 3.1 Compliant Intervention Interface
Correction data is collected by human demonstrators to rectify policy failures. Unlike initial demonstrations that establish baseline behaviors, correction data specifically targets failure modes observed during policy deployment. Correction data is most effective when it corrects failures in policy-induced state distributions [ross2011reduction]. The interface through which these corrections are collected significantly impacts the quality of correction data, which should be intuitive for demonstrators, capture critical corrective information at precise moments of failures, and facilitates collecting data close to the base policy state-action distribution.
There are two types of correction collection methods: Off-policy correction is when humans observe failures of the base policy during deployment, and then recollect extra offline demonstrations to address failure cases. This approach is most commonly used for improving Behavior Cloning policy performance due to its simplicity - it requires no additional infrastructure beyond the original data collection setup. However, the resulting demonstrations may fail to cover all the failure cases or deviate from the original state-action distribution. We focus on on-policy correction instead, where humans can monitor policy execution and intervene on the spot when failures occur or are anticipated. This approach allows humans to provide corrections more targeted to the base policy’s failure cases. However, challenges still exist for an intervention system:
@@
  • Distribution shift. The human-intervened state-action distribution may deviate significantly from the original distribution. Additionally, the non-smooth transition above could bring in disturbances and add to the distribution shift.
  • Indirect correction brings errors. Correction is commonly implemented via teleoperation interfaces such as spacemouse or joysticks [he2025uncertainty, chen2025conrft]. With spatial mismatch and teleoperation latency, it is hard for the demonstrator to instantly provide accurate corrections upon intervention starts without going through a short adjustment period.
@@
  • Delta correction instead of take-over correction. Unlike take-over correction, where the demonstrator has no idea of the policy’s original intention once taking over, we propose a novel on-policy delta correction method: we let the robot policy executes continuously while the human applies forces to the robot with a handle mounted on the end effector, resulting in delta actions on top of the policy action. The human demonstrator can always sense the policy’s intention through haptic feedback, and easily control the magnitude of intervention by the amount of force applied to the handle. As a result, delta correction ensures smooth intervention data and limits the human from providing very large corrections. The approach is also intuitive as the human can directly move the robot towards desired correction directions.
  • Correction interface with compliance control. In order to apply delta correction over a running policy, we provide a compliant interface that allows humans to safely intervene and apply force to the robot to affect its behaviors at any time, as shown in Fig. 2. We design a kinesthetic correction hardware setup with a detachable handle for human to hold when correcting, and allows easy tool-swapping for different tasks. We run a compliance controller (specifically admittance control) in the background to respond to both contact forces and human correction forces, allowing the human to influence but not completely override the policy execution. The admittance controller uses a constant stiffness $\sim$$1000\text{\,}\mathrm{N}\mathrm{/}\mathrm{m}$ to allow easy human intervention and ensure accurate tracking.
@@
Figure 3: Policy Update Methods. Left: Common policy update methods - retraining and finetuning. Right: Ours. The base policy runs at $1\text{\,}\mathrm{H}\mathrm{z}$. It takes in images $I_{t}$ and proprioceptions $P_{t}$ and predicts 32 frames of end-effector poses $A_{t}^{p}=\{A_{t_{0}}^{p},A_{t_{1}}^{p},...,A_{t_{T_{a}}}^{p}\}$ spaced $0.1\text{\,}\mathrm{S}\mathrm{e}\mathrm{c}\mathrm{o}\mathrm{n}\mathrm{d}\mathrm{s}$ apart. The Compliant Residual Policy runs at $50\text{\,}\mathrm{H}\mathrm{z}$. It takes in additional force inputs $F_{t}$ and predicts 5 frames of delta poses $\Delta A_{t}^{p}=\{\Delta A_{t_{0}}^{p},\Delta A_{t_{0}+1}^{p},...,\Delta A_{t_{0}+{T_{r}}}^{p}\}$ and target forces $A_{t}^{f}$ spaced $0.02\text{\,}\mathrm{S}\mathrm{e}\mathrm{c}\mathrm{o}\mathrm{n}\mathrm{d}\mathrm{s}$ apart. The combined poses of $A_{t}^{p}$ and $\Delta A_{t}^{p}$, and target forces $A_{t}^{f}$ are taken by an admittance controller to command the robot.
@@ 3.2 Compliant Residual Policy
Compliant residual policy formulation. Our policy directly learns corrective behavior from the human delta correction data, as shown in Fig. 3. It takes as input the same visual and proprioceptive feedback as the base policy but with a shorter horizon. It also takes in an extra force modality, which is available using our Compliant Intervention Interface. The policy outputs five frames of actions at a time, corresponding to $0.1\text{\,}\mathrm{s}$ of execution time when running at $50\text{\,}\mathrm{H}\mathrm{z}$. The action space is 15-dimensional: the first nine dimensions represent the SE3 delta pose from the base policy action to the robot pose command [chi2023diffusion], while the later six dimensions represent the expected wrench (force and torque) the robot should feel from external contacts. Both the robot pose command and the expected wrench are sent to a standard admittance controller for execution with compliance.
The residual policy directly uses the base policy’s frozen image encoder [radford2021learning, amir2021deep, xu2023jacobinerf] to extract an image embedding, a temporal convolution network [van2016wavenet] to encode the force vectors, followed by fully-connected layers to decode actions.
Advantages. This formulation provides the following advantages:
@@
Training strategy. In prior work, a residual policy is trained either in simulation with RL [ankile2024imitation, yuan2024policy] to give it sufficient coverage of the state distribution, or in the real world with pre-collected behavior cloning data [luo2024precise]. In this work, we train the Compliant Residual Policy completely on the small amount of new real-world correction data with the following strategies:
@@
  • Prioritize correction data over no-correction (zero residual action) data. Similar to [liu2022robot], we alter the data sample frequency during training based on whether they have human correction or not. Specifically, since the moment of correction start indicates where the current policy performs badly followed by immediate action to fix it, we sample data more frequently for a short period immediately after correction starts. Our real-world ablations (§ 4.5) demonstrate that our training strategies improve the quality of the residual policy and the overall success rate.
@@ 4 Evaluation
[…]
@@ 4.1 Contact-Rich Manipulation Tasks
Figure 5: Belt Assembly Task. (a) Policy rollout of [Compliant Residual] policy trained with [On-Policy Delta] data, demonstrating accurate force-position coordination and adaptation. (b) Different test scenarios. (c) Typical failure cases of the base policy: missing the slot by going too high above the pulley; tilting the belt and causing it to slip off the pulley.
Book Flipping: As shown in Fig. 4 (a), this task is to flip up books on a shelf. Starting with one or more books lying flat on the shelf, the robot first insert fingers below the book, then rotate the book up and push them firmly against the shelf to let them stand on their own. The base policy is trained with 150 demonstrations.
This task is challenging as it involves rich use of physical contacts and forceful strategies [hou2020manipulation]. A position-only strategy always fails immediately by triggering large forces, so we execute all policies through the same admittance controller. The task success requires high precision in both motion and force to accurately align the fingers with the gap upon insertion, and to provide enough force to rotate heavy books and make the books stand firmly.
@@
Belt Assembly: As shown in Fig. 5 (a), this task is to assemble a thin belt onto two pulleys, which is part of the NIST board assembly challenge [kimble2020benchmarking]. Starting with the belt grasped by the gripper, the robot needs to first thread the belt over the small pulley, next move down while stretching the belt to thread its other side on the big pulley, then rotate $180\text{\,}\mathrm{\SIUnitSymbolDegree}$ around the big pulley to tighten the belt, and finally pull up to release the belt from the gripper. The base policy is trained with 405 demonstrations.
The task is challenging as it requires both position and force accuracy throughout the process. Specifically, the belt is thin and soft so the initial alignments onto the pulleys are visually ambiguous. Also, since the belt is not stretchable, there is more resistance force and less position tolerance as the belt approaches the second pulley, requiring a policy with good force-position coordination and adaptation. We ran 32 rollouts across four different initial board positions and four grasp locations (Fig. 5 (b)).
@@ 4.2 Base Policy and its Failure Modes
We trained a diffusion policy [chi2023diffusion] that takes in past images from a wrist-mounted camera and robot proprioception observations, and predicts a future position-based action trajectory. To isolate the contribution of force inputs versus human corrections, we trained diffusion policies with and without force inputs as baselines for the belt assembly task.
@@
The belt assembly base policy achieves a 15.6% success rate. Adding force input increases the base policy success rate to 43.8%. Common failure cases include (Fig. 5 (c)): (1) Missed slotting: the fingertip goes too high or too low, causing the belt to miss the slot on the big pulley. (2) Belt slippage: the fingers pull the belt in the wrong direction, causing the belt to tilt and slip off the pulley.